OpenStudy (anonymous):

Simplify. (will post everything, what to simplify and answer choices inside!) Thanks! :) please explain :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Answer choices A,B,C,D from top to bottom :) thanks

OpenStudy (anonymous):

\[\sqrt{64(-1)}=\sqrt{-1}\sqrt{64}\]

OpenStudy (anonymous):

\[i8\]

OpenStudy (anonymous):

so 8i ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cuz square root of -1 is i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and that makes my answer C. 8i right??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

awesome :) thanks!